# Compressed data structures for Heegaard splittings

### v1

@@ 3. Data structure for Heegaard diagrams
Attach $f^{k_{i,1}}_{i,1}$ to $w_{\gamma}$. Call $(t^{\prime})_{i}^{1}$ the triangle adjacent to $t_{i}$ by the edge $f_{i,2}$; note that $(t^{\prime})_{i}^{1}$ is also incident to $v_{i}$. Again, because of the barycentric subdivision, there is exactly one edge $(f^{\prime})_{i,2}^{1}$ distinct from $(f^{\prime})_{i,1}^{1}=f_{i,2}$ incident to $v_{i}$. The pair $((f^{\prime})^{1}_{i,1},(f^{\prime})^{1}_{i,2})$ defines an elementary arc in $(t^{\prime})_{i}^{1}$ given by an intersection sequence $((f^{\prime})^{1}_{i,1},(f^{\prime})^{1}_{i,2})$ (we here hide the indices of intersection to simplify the notation), see the dashed purple lines in Figure 3(a). Similarly and recursively, the triangle $(t^{\prime})^{2}_{i}$ that shares $(f^{\prime})^{1}_{i,2}$ with $(t^{\prime})^{1}_{i}$ has exactly one other edge, call it $(f^{\prime})^{2}_{i,2}$, incident to $v_{i}$. Because $f^{k_{i,1}}_{i,1}$ is incident to $v_{i}$, we can repeat the process until $(f^{\prime})^{n}_{i+1,2}=f^{k_{i,1}}_{i,1}$ for some $n\leq 3|T^{\prime}|$. We then append $(f^{\prime})^{1}_{i,1},(f^{\prime})^{2}_{i,1},\dots,(f^{\prime})^{n}_{i,1},f_{i+1,1}$ to $w_{\gamma}$. Since the curve $\gamma$ is assumed simple, no edge or vertex occurs twice in the list $E_{T}(\gamma)$, meaning that this process takes time, at most, $O(|T^{\prime}|)=O(|T|)$. We then remove edges in $T^{\prime}-T$ from $w_{\gamma}$ and cyclically reduce it to guarantee that it represents a normal curve. Finally, $w_{\gamma}$ can be transformed into a trivial SLP, $I_{T}(\gamma)$, in time $O(|w_{\gamma}|)=O(|T|)$.
Now, assume that #$\#\gamma\geq 2$ and suppose we separately apply the above procedure to each component $c_{1},\dots,c_{\#\gamma}$ of $\gamma$, obtaining uncompressed sequences $w_{1},\dots,w_{\#\gamma}$. Although these sequences individually represent curves that are isotopic to the components of $\gamma$, tracing them might cause some intersection between two components, say $c_{1}$ and $c_{2}$, if $E_{T}(c_{1})$ and $E_{T}(c_{2})$ shared a vertex, as shown in Figure 3(b). One can detect these cases in time $O(\#\gamma\times\|E_{T}(\gamma)\|)$ and solve the intersections by exchanging the parts of the sequence $w_{1}$ with the parts of the sequence $w_{2}$ that cobound bigons, ultimately performing the surgery in the bottom of Figure 3(b). We then cyclically reduce each curve $w_{1},\dots,w_{\#\gamma}$—note that the cyclic reductions cannot form new bigons between components—and compute the trivial SLPs. ∎
@@
For the second part, note that while $I_{T}(\beta)$ is, by construction, standard in $T$, it might not be normal. We can use the main algorithm of [42] to deterministic compute, in time $O(\|I_{T}(\beta)\|^{10})$, cyclic reduced SLPs for $\beta$ with complexity $O(\|I_{T}(\beta)\|^{2})$. Alternatively, this can done using a randomized algorithm in time $O(\|I_{T}(\beta)\|^{3})$ with some probability of error at most $\delta$ and in a unit-cost RAM model of bit-size $O(\|I_{T}(\beta)\|+\lceil 1/\delta\rceil)$. Part (c) of Lemma 2.2 finishes the proof. ∎
@@ 4. Algorithms on Heegaard diagrams
We here describe some polynomial time algorithms and operations on a genus $g$ Heegaard diagram $(T,E_{T}(\alpha),N_{T}(\beta))$ with complexity $m$ and representing a closed 3-manifold $M$. For convenience, the algorithms and their complexity are summarized in Table 1. For the reported complexities, we assume a unit-cost RAM model of bit-size $\log|\beta\cap E|=O(\|N_{T}(\beta)\|/|T|)$.
@@
Use a generalization of Theorem 3.3 [17][Theorem 6.5] to trace the street complex $S(T,\beta)$, this naturally induces triangulations of the components in $T\backslash\beta$, which, following [18], we call pieces. We compute the Euler characteristic of each piece in time $O(|T|)$ [17], ultimately giving their genus. The vectors $N_{T}(\beta)$ define a Heegaard splitting if and only if the only piece is a $2g$-punctured sphere. ∎
@@
An algorithm for this problem is described by Lackenby in Theorem 6.3 of [34], where we implicitly use the content of his Corollary 6.5, that is, the number of bigons between the $\alpha$ and $\beta$ curves is upper bounded by $154|T|$. ∎
@@
There is an algorithm to check, in time $O(\operatorname{poly}(m,|T|))$, whether two Heegaard diagrams, $\beta$ and $\beta^{\prime}$ of complexity at most $m$ are isotopic.
@@
Once more, this follows from Lackenby’s paper [34], this time, from Theorem 1.2. ∎
@@
DO A DEHN TWIST Compute the action of a Dehn tiwst about a normal curve $s$ with complexity $m^{\prime}$ on $\beta$ $O(((m+m^{\prime})|T|)^{10})$ Theorem 4.9
GET $\pi_{1}$ Compute a presentation of the fundamental group of $M$, with relations given by SLPs $O(m|T|)$ Theorem 4.10
@@
GET $\beta_{1}$ Compute the first Betti number of $M$ $\widetilde{O}(mg^{3}+m|T|)$ Theorem 4.13
@@
Refer to the two moves of Figure 4. These modifications of the underlying are known as Pachner moves and have received special attention in low-dimensional topology because any two triangulations are related by finitely many applications of them [45]. Here, we assume that no such a move deletes an edge in $E_{T}(\alpha)$, otherwise, we would not have a marked triangulation as an output. Move $(3,1)\to(1,3)$ corresponds to the simple deletion of the relevant components of $N_{T}(\beta)$, move $(1,3)\to(3,1)$ involves adding new components whose values are corner coordinates. Both are implemented in time $O(1)$ in our data structure for $T$. Similarly, moves $(2,2)\to(2,2)$ can be implemented in time $O(1)$ directly from computing the corner coordinates and updating the triangulation. ∎
@@
Compute the street complex $S(T,\beta)$, recall that it has complexity at most $O(|T|)$. Using breadth-first in the dual graph of $S(T,\beta)$, find, in time $O(|T|^{2})$, the shortest arc $p$ that connects faces adjacent to $b_{1}$ and $b_{2}$ and that does not cross any component of $\beta$, which can be avoided by adding infinite weights to the edges of $S(T,\beta)$ not in $T$. Note that such an arc must exist because no component of $\beta$ can be separating in the surface. Moreover, because $p$ is shortest and therefore intersects at most $O(|T|)$ streets in $S(T,\beta)$, we can compute using part (b) of Theorem 3.3, an SLP of complexity $O(m|T|+|T|)=O(m|T|)$ of an intersection sequence of $p$ with respect to $T$, call it $I_{T}(p)$.
Suppose that $p$ is not trivial, that is, it is not only a single point. By Theorem 6.2 of [18], find, in time $O(m|T|)$, SLPs of the intersection sequences of $b_{1}$ and $b_{2}$ with respect to $T$ starting at edges adjacent to the endpoint faces of $p$, call them $I_{T}(b_{1})$ and $I_{T}(b_{2})$. Finally, let $I_{T}(b)=I_{T}(b_{2})^{-1}\circ I_{T}(p)^{-1}\circ I_{T}(b_{1})\circ I_{T}(p)$. Note that $I_{T}(b)$ already is an intersection sequence of a normal curve in $T$. For, as $p$ is a shortest path, it is normal in $S(T,\beta)$, which implies it is also normal in $T$. Moreover, $p$ has as endpoints junctions in $S(T,\beta)$ as, if it had steets as endpoints, there would be a shorter arc connecting $b_{1}$ to $b_{2}$ starting at one of the ends of the street. Therefore, any other reduction in $I_{T}(b)$ would necessarily arise by $p$ entering a junction of $S(T,\beta)$ through a non-redundant port $e$ that is also intersected by either $b_{1}$ or $b_{2}$. But then, the other triangle in $T$ adjacent to $e$ is adjacent to either $b_{1}$ or $b_{2}$, implying that $p$ is not shortest. Part (c) of Lemma 2.2 finishes this case.
If $p$ is trivial, then it means that $b_{1}$ and $b_{2}$ bound a same face in $S(T,\beta)$. Nonetheless, because the curves are not to isotopic, there exists at least one face bounded by $b_{1}$ and not bounded by $b_{2}$. We call such a face good. Choose, in time $O(|T|)$, one non-good face $t_{1}$ that is adjacent to a good face through an edge $e_{1}$, see Figure 5. Note that $e_{1}$ is the endpoint of a street $s$ of sides in $b_{1}$ and $b_{2}$. Look for the other end of this street, $e_{2}$—again, because $b_{1}$ and $b_{2}$ are not isotopic, the street has another end. Let $t_{2}$ be the face (in fact, a junction) not equal to $s$, but also adjacent to $e_{2}$. Compute, once again through [18][Theorem 6.2], the sequence of faces adjacent to $b_{1}$ starting at $t_{1}$ and going all the way up to $t_{2}$ without crossing $s$, call this sequence $I_{T}(\widetilde{b_{1}})$. Compute the related sequence for $b_{2}$, this time stating at $t_{2}$, call it $I_{T}(\widetilde{b_{2}})$. Let $I_{T}(b)=I_{T}(\widetilde{b_{2}})\circ I_{T}(\widetilde{b_{1}})$, note that it is also normal in $T$ as $b_{1}$ and $b_{2}$ are normal and it enters $t_{1}$ and $t_{2}$ through distinct edges. ∎
Figure 5. Left: diagram for the proof of Theorem 4.5. The red and blue curves represent $b_{1}$ and $b_{2}$, the black lines are other edges of $S(T,\beta)$. The faces $s$, $t_{1}$, and $t_{2}$ are also indicated. Right: diagram for the proof of Theorem 4.6. Figure inspired by [16].
@@
A similar algorithm for uncompressed intersection sequences as inputs is described in [16]; here we extend it to our data structure. First, let $\gamma$ be a normal (multi)curve with respect to a triangulation $T$. Having fixed a triangle in $T$, we call its corner coordinates the number of elementary arcs of each type. In particular, note that for the triangle $t$ with sides $e$, $f$, and $g$, the number of elementary arcs between sides $e$ and $f$ is $1/2\times[e(\gamma)+f(\gamma)-g(\gamma))]$ and symmetrically for the other cases. Given $e(\gamma),f(\gamma)$, and $g(\gamma)$, one can use these formulas to compute the corner coordinates in $t$ in time $O(1)$.
Choose any such face $t$ and modify it by adding a triangle $t_{1}$ of sides $e_{1},f_{1},g_{1}$ to its center, together with other 6 edges $e^{(1)}_{1},e^{(2)}_{1},f^{(1)}_{1},f^{(2)}_{1},g^{(1)}_{1},g^{(2)}_{1}$ connecting the vertices of $t_{1}$ to the vertices of $t$, see Figure 5. Define the triangulation $T_{1}$ as $T$ with $t_{1}$ deleted and update the normal coordinates of $\beta$ to be a vector in $\mathbb{N}^{|T|+9}$, where the new components are set to
@@
Now let $T_{2}$ be a fixed triangulation of the punctured torus with a triangular boundary component $t_{2}$, endowed with an edged curve $a$ and a normal curve $b$, both disjoint from $t_{2}$. Glue $T_{1}$ to $T_{2}$ by identifying $t_{1}$ with $t_{2}$. This forms a triangulation of $\Sigma_{g+1}$, with $\alpha\cup\{a\}$ being an edged multicurve and $\beta\cup\{b\}$ naturally giving a Heegaard diagram of a stabilization of $(\Sigma_{g},\alpha,\beta)$. ∎
@@
For each $a\in\alpha$, check the normal vectors of $N_{T}(\beta)$, in time $O(g\|E_{T}(\alpha)\|)$, for a $a$ that intersects only a single component of $\beta$ and only once. Suppose that there is such an $a$ and that $a,b$ is the detected trivial stabilization pair. In time $O(\|E_{T}(a)\|)=O(|T|)$, cut $T$ at $a$ and attach disks at the holes to have a triangulation of the surface of genus $g-1$. ∎
@@
We divide the algorithm into two parts, corresponding to the two hypotheses of a reducibility of diagrams. For the first, apply [34][Theorem 6.3] to find, in polynomial time on the input, an isotopic diagram $(\alpha,\beta^{\prime})$ in efficient position where the complexity of $\beta$ increases to $m^{\prime}=m(1+2|T|)$. We note that although this subroutine is polynomial time on the input, the author of [34] does not provide an estimation of the polynomial degree and any attempt on our part to give one would be out of the scope of this paper. Because we made the diagram efficient, there can only be a component $a\in\alpha$ isotopic to some $b^{\prime}\in\beta^{\prime}$ if $e(b^{\prime})=0$ for all $e\in E_{T}(\alpha)$. If that is the case, we trace $S(T,b^{\prime})$. Since $b^{\prime}\cap a=\emptyset$, the curve $a$ is still fully in the edges of the street complex. Nonetheless, $a$ is isotopic to $b^{\prime}$ if and only if they cobound a cylinder [19], a case that can be determined in time $O(\|S(T,b^{\prime})\|)$. Checking this for every pair of curves $a$ and $b^{\prime}$ that do not intersect gives a test for the first hypothesis.
If the above test returns no clear reduction pair, we look for a separating curve disjoint from $\alpha$ and $\beta$. Trace $S(T,\beta)$ and add infinite weight to the edges of the dual graph corresponding to crossings of $E_{S(T,\beta)}(\beta)$, therefore defining a cellular embedding for the $2g$-punctured sphere $\Sigma_{g}\backslash\beta$. In addition, for each street in $E_{S(T,\beta)}(\beta)$, we can determine in time $O(\|E_{T}(\alpha)\|\times m|T|)$ with dynamic programming if there is an occurrence of an edge in $E_{T}(\alpha)$ in the crossing sequence. Whenever that is the case, we also add infinite weights to the edges of the dual graph of $E_{S(T,\beta)}(\beta)$ that cross the non-redundant ports bounding the street. Because each street is either a disk or a segment of an edge of $T$, adding the infinite weights defines a cellular embedding $T^{\prime}$ of a manifold homeomorphic to $\Sigma^{\prime}=\Sigma_{g}\backslash(\alpha\cup\beta)$. We note that $\Sigma^{\prime}$ is a union of (potentially one) punctured spheres and that an essential separating curve exists if and only if all components are not disks. Assume that that is the case, then using the techniques of [17]—refer to Section 2.3 of [9] for details—one can compute, in time $O(|T^{\prime}|)$ an essential curve $s$ in $\Sigma_{g}\backslash(\alpha\cup\beta)$ expressed as an SLP of complexity $O(m|T|+|T|)=O(m|T|)$. ∎
@@
Given a normal curve $s$ in $T$ of complexity $\|N_{T}(s)\|=m^{\prime}$, there is an algorithm to compute, in time $O(((m+m^{\prime})|T|)^{10})$, the diagram $(T,E_{T}(\alpha),N_{T}(\tau_{s}(\beta))$.
@@
This algorithm is described by [49, 55]. If one requires only the intersection sequences $I_{T}(\tau_{s}(\beta))$, the algorithm can be executed in linear time on the inputs, but the output might just be standard with respect to the triangulation. Therefore, as in the proof of Theorem 3.6, we are forced to use the subroutines described by the same authors to cyclic reduce—either in time $O(((m+m^{\prime})|T|)^{10})$ with a deterministic algorithm or in time $O(((m+m^{\prime})|T|)^{3})$ with a randomized one—the multicurve given by $I_{T}(\tau_{s}(\beta))$. ∎
@@
A presentation of $\pi_{1}(M)$ with the relations given by SLPs can be computed in time $O(m|T|)$.

### v2

@@ 3. Data structure for Heegaard diagrams
Attach $f^{k_{i,1}}_{i,1}$ to $w_{\gamma}$. Call $(t^{\prime})_{i}^{1}$ the triangle adjacent to $t_{i}$ by the edge $f_{i,2}$; note that $(t^{\prime})_{i}^{1}$ is also incident to $v_{i}$. Again, because of the barycentric subdivision, there is exactly one edge $(f^{\prime})_{i,2}^{1}$ distinct from $(f^{\prime})_{i,1}^{1}=f_{i,2}$ incident to $v_{i}$. The pair $((f^{\prime})^{1}_{i,1},(f^{\prime})^{1}_{i,2})$ defines an elementary arc in $(t^{\prime})_{i}^{1}$ given by an intersection sequence $((f^{\prime})^{1}_{i,1},(f^{\prime})^{1}_{i,2})$ (we here omit the intersection indices to simplify the notation), see the dashed purple lines in Figure 5(a). Similarly and recursively, the triangle $(t^{\prime})^{2}_{i}$ that shares $(f^{\prime})^{1}_{i,2}$ with $(t^{\prime})^{1}_{i}$ has exactly one other edge incident to $v_{i}$, call it $(f^{\prime})^{2}_{i,2}$. Because $f^{k_{i,1}}_{i,1}$ is incident to $v_{i}$, we can repeat the process until $(f^{\prime})^{n}_{i+1,2}=f^{k_{i,1}}_{i,1}$ for some $n\leq 3|T^{\prime}|$. We then append $(f^{\prime})^{1}_{i,1},(f^{\prime})^{2}_{i,1},\dots,(f^{\prime})^{n}_{i,1},f_{i+1,1}$ to $w_{\gamma}$. Since the curve $\gamma$ is assumed simple, no edge or vertex occurs twice in the list $E_{T}(\gamma)$, implying that this process takes time $O(|T^{\prime}|)=O(|T|)$. We then remove edges in $T^{\prime}-T$ from $w_{\gamma}$ and cyclically reduce it to ensure that it represents a normal curve. Finally, $w_{\gamma}$ can be transformed into a trivial SLP, $I_{T}(\gamma)$, in time $O(|w_{\gamma}|)=O(|T|)$.
Now, assume that $\gamma\geq 2$ and suppose we separately apply the above procedure to each component $c_{1},\dots,c_{\#\gamma}$ of $\gamma$, obtaining uncompressed sequences $w_{1},\dots,w_{\#\gamma}$. Although these sequences individually represent curves that are isotopic to the components of $\gamma$, tracing them might create intersections between two components, say $c_{1}$ and $c_{2}$, if $E_{T}(c_{1})$ and $E_{T}(c_{2})$ shared a vertex, as shown in Figure 5(b). One can detect these cases in time $O(\#\gamma\times\|E_{T}(\gamma)\|)$ and solve the intersections by exchanging the parts of the sequence $w_{1}$ with the parts of the sequence $w_{2}$ that cobound bigons, ultimately performing the surgery in the bottom of Figure 5(b). We once more cyclically reduce each curve $w_{1},\dots,w_{\#\gamma}$ (note that the cyclic reductions cannot form new bigons between components) and compute the trivial SLPs. ∎
@@
For the second part, note that while $I_{T}(\beta)$ is, by construction, standard in $T$, it might not be normal. We can use the main algorithm of [miyazaki1997improved] to deterministic compute, in time $O(\|I_{T}(\beta)\|^{10})$, cyclic reduced SLPs for $\beta$ with complexity $O(\|I_{T}(\beta)\|^{2})$. Alternatively, this can be done using a randomized algorithm in time $O(\|I_{T}(\beta)\|^{3})$ with some probability of error at most $\delta$ and in a unit-cost RAM model of bit-size $O(\|I_{T}(\beta)\|+\lceil 1/\delta\rceil)$. Part (c) of Lemma 2.3 finishes the proof. ∎
@@ 4. Algorithms on Heegaard diagrams
We here describe some polynomial time algorithms and operations on a genus $g$ Heegaard diagram $(T,E_{T}(\alpha),N_{T}(\beta))$ with complexity $m$ and representing a closed 3-manifold $M$. For convenience, the algorithms and their complexity are summarized in Table 1. For the reported times, we assume a unit-cost RAM model of bit-size $\log|\beta\cap E|=O(\|N_{T}(\beta)\|/|T|)$.
@@
Use a generalization of Theorem 3.2 [erickson2002optimally][Theorem 6.5] to trace the street complex $S(T,\beta)$, this naturally induces triangulations of the components in $T\backslash\beta$, which, following [Erickson2013], we call pieces. We compute the Euler characteristic of each piece in time $O(|T|)$ [erickson2002optimally], ultimately giving their topological type. The vectors $N_{T}(\beta)$ define a Heegaard splitting if and only if the only piece is a $2g$ punctured sphere. ∎
@@
Lackenby describes an algorithm for this problem in Theorem 6.3 of [lackenby2024some], where we implicitly use the content of his Corollary 6.5, namely that the number of bigons between the $\alpha$ and $\beta$ curves is upper bounded by $154|T|$. ∎
@@
There is an algorithm to check, in time $O(\operatorname{poly}(m,|T|))$, whether two Heegaard diagrams, $\beta$ and $\beta^{\prime}$, of complexity at most $m$, are isotopic.
@@
Once more, this follows from Lackenby’s paper [lackenby2024some], this time, from Theorem 1.2. ∎
@@
DO A DEHN TWIST Compute the action of a Dehn twist about a normal curve $s$ with complexity $m^{\prime}$ on $\beta$ $O(((m+m^{\prime})|T|)^{10})$ Theorem 4.9
GET $\pi_{1}$ Compute a presentation of the fundamental group of $M$ $O(m|T|)$ Theorem 4.10
@@
First, let $\gamma$ be a normal (multi)curve with respect to a triangulation $T$. Having fixed a triangle in $T$, we call its corner coordinates the number of elementary arcs of each type. In particular, note that for the triangle $t$ with sides $e$, $f$, and $g$, the number of elementary arcs between sides $e$ and $f$ is $1/2\times[e(\gamma)+f(\gamma)-g(\gamma))]$ and symmetrically for the other cases. Given $e(\gamma),f(\gamma)$, and $g(\gamma)$, one can use these formulas to compute the corner coordinates of $t$ in time $O(1)$.
Now refer to the two moves of Figure 6. These modifications of the underlying triangulation are known as Pachner moves and have received special attention in low-dimensional topology because any two triangulations are related by finitely many applications of them [pachner1991pl]. Here, we assume that no such move deletes an edge in $E_{T}(\alpha)$; otherwise, we would not have a marked triangulation as an output. Move $(3,1)\rightarrow(1,3)$ corresponds to the simple deletion of the relevant components of $N_{T}(\beta)$, move $(1,3)\rightarrow(3,1)$ involves adding new components whose values are corner coordinates. Both are implemented in time $O(1)$ in our data structure for $T$. Similarly, moves $(2,2)\rightarrow(2,2)$ can be implemented in time $O(1)$ directly by computing the corner coordinates and updating the triangulation. ∎
@@
Compute the street complex $S(T,\beta)$; recall that it has complexity of $O(|T|)$. Using breadth-first in the dual graph of $S(T,\beta)$, find, in time $O(|T|^{2})$, the shortest arc $p$ that connects faces adjacent to $b_{1}$ and $b_{2}$ and that does not cross any component of $\beta$, which can be avoided by adding infinite weights to the edges of $S(T,\beta)$ not in $T$. Note that such an arc must exist because no component of $\beta$ can be separating in the surface. Moreover, because $p$ is shortest and, therefore, intersects at most $O(|T|)$ streets in $S(T,\beta)$, we can compute, using Part (b) of Theorem 3.2, an SLP of complexity $O(m|T|+|T|)=O(m|T|)$ of an intersection sequence of $p$ with respect to $T$, call it $I_{T}(p)$.
Suppose that $p$ is not trivial, that is, it is not just a single point. By Theorem 6.2 of [Erickson2013], find, in time $O(m|T|)$, SLPs of the intersection sequences of $b_{1}$ and $b_{2}$ with respect to $T$ starting at edges adjacent to the endpoint faces of $p$, call them $I_{T}(b_{1})$ and $I_{T}(b_{2})$. Finally, let $I_{T}(b)=I_{T}(b_{2})^{-1}\circ I_{T}(p)^{-1}\circ I_{T}(b_{1})\circ I_{T}(p)$. Note that $I_{T}(b)$ already is an intersection sequence of a normal curve in $T$. For, since $p$ is a shortest path, it is normal in $S(T,\beta)$, which implies that it is also normal in $T$. Moreover, $p$ has junctions as endpoints; if it had streets as endpoints, there would be a shorter arc connecting $b_{1}$ to $b_{2}$ starting at one of the ends of the street. Therefore, any other reduction in $I_{T}(b)$ would necessarily arise by $p$ entering a junction of $S(T,\beta)$ through a non-redundant port $e$ that is also intersected by either $b_{1}$ or $b_{2}$. But then, the other triangle in $T$ adjacent to $e$ is adjacent to either $b_{1}$ or $b_{2}$, implying that $p$ is not shortest. By repeated applications of Part (c) of Lemma 2.3, one can compute, in time $O(|T|\times m)$, the number of (unsigned) occurrences of each edge of $E$ in the SLP of $I_{T}(b)$, resulting in $N_{T}(b)$.
[…]
Figure 7. Left: diagram for the proof of Theorem 4.5. The red and blue curves represent $b_{1}$ and $b_{2}$, the black lines are the other edges of $S(T,\beta)$. The faces $s$, $t_{1}$, and $t_{2}$ are also indicated. Right: diagram for the proof of Theorem 4.6. Figure inspired by [ennes2025hardnesscomputationquantuminvariants].
@@
A similar algorithm for (uncompressed) intersection sequences as input is described in [ennes2025hardnesscomputationquantuminvariants]; here we extend it to our data structure. Choose any face $t$ of $T$ and modify it by adding a triangle $t_{1}$ of sides $e_{1},f_{1},g_{1}$ to its center, together with 6 other edges $e^{(1)}_{1},e^{(2)}_{1},f^{(1)}_{1},f^{(2)}_{1},g^{(1)}_{1},g^{(2)}_{1}$ connecting the vertices of $t_{1}$ to the vertices of $t$, see Figure 7. Define the triangulation $T_{1}$ as $T$ with $t_{1}$ deleted and update the normal coordinates of $\beta$ to be a vector in $\mathbb{N}^{|T|+9}$, where the new components are set to
@@
Now let $T_{2}$ be a fixed triangulation of the punctured torus with a triangular boundary component $t_{2}$, endowed with an edged curve $a$ and a normal curve $b$, both disjoint from $t_{2}$. Glue $T_{1}$ to $T_{2}$ by identifying $t_{1}$ with $t_{2}$. This forms a triangulation of $\Sigma_{g+1}$, with $\alpha\cup\{a\}$ and $\beta\cup\{b\}$ naturally giving a Heegaard diagram of a stabilization of $(\Sigma_{g},\alpha,\beta)$. ∎
@@
For each $a\in\alpha$, check the normal vectors of $N_{T}(\beta)$, in time $O(g\|E_{T}(\alpha)\|)$, for an $a$ that intersects only a single component of $\beta$ and only once. Suppose that there is such an $a$ and that $a,b$ is the detected trivial stabilization pair. In time $O(\|E_{T}(a)\|)=O(|T|)$, cut $T$ at $a$ and attach disks to the holes, therefore giving a triangulation of the surface of genus g1g-1. ∎
@@
We divide the algorithm into two parts, corresponding to the two hypotheses of reducibility of diagrams, namely the existence of an isotopic pair of meridians $a\in\alpha$ and $b\in\beta$, or of an essential separating curve that splits $\Sigma_{g}$ into subdiagrams. For the first case, apply [lackenby2024some][Theorem 6.3] to find, in polynomial-time on the input, an isotopic diagram $(\alpha,\beta^{\prime})$ in efficient position. This procedure increases the complexity of $\beta$ to $m^{\prime}=m(1+2|T|)$ We note that although this subroutine is polynomial-time on the input, the author of [lackenby2024some] does not provide an estimation of the polynomial degree, and any attempt on our part to give one would be out of the scope of this paper. Because we made the diagram efficient, there can only be a component $a\in\alpha$ isotopic to some $b^{\prime}\in\beta^{\prime}$ if $e(b^{\prime})=0$ for all $e\in E_{T}(\alpha)$. If that is the case, we trace $S(T,b^{\prime})$. Since $b^{\prime}\cap a=\emptyset$, the curve $a$ is still fully in the edges of the street complex. Nonetheless, $a$ is isotopic to $b^{\prime}$ if and only if they cobound a cylinder [farb2011primer], a case that can be determined in time $O(|S(T,b^{\prime})|)$. Checking this for every pair of curves $a$ and $b^{\prime}$ that do not intersect—there can be at most $g^{2}$ of these pairs—gives a test for the first hypothesis.
Figure 8. Diagram illustrating the proof of Theorem 4.8. Left: part of the overlay graph defined by the $\beta$ curves (blue) with respect to a triangulation (black) and the corresponding embeddings of a segment of the $\alpha$ curves (red). Redundant ports are shown in dashed. A cellular embedding of the surface $\Sigma^{\prime\prime}$ can be obtained by puncturing the diagram along both blue and red curves. Right: another cellular embedding for $\Sigma^{\prime\prime}$ in which we enlarge the holes punctured when cutting $\alpha$ by moving redundant ports in $E_{T}(\alpha)$ to both ends of the streets.
If the above test returns no clear reduction pair, we look for a separating curve, $s$, disjoint from $\alpha$ and $\beta$, which exists if and only if $s$ is non-contractable in $\Sigma^{\prime\prime}=\Sigma_{g}\backslash(\alpha\cup\beta)$. Because $\Sigma^{\prime}=\Sigma_{g}\backslash\beta$ is a $2g$ puncture sphere, $\Sigma^{\prime\prime}$ is a union of (potentially one) punctured spheres, and it embeds a non-contractable curve if and only if one of its components is not a disk. This means that we only need to compute a cellular embedding of $\Sigma^{\prime\prime}$ and check its topological type. We start computing a triangulation for $\Sigma^{\prime}$ as in the proof of Theorem 4.5 by tracing $S(T,\beta)$ and disconnecting from the gluing rules the edges of $E_{S(T,\beta)}(\beta)$. We then want to further cut $\Sigma^{\prime}$ along $\alpha$. But, because $\alpha$ is edged with respect to $T$, each segment of $E_{T}(\alpha)$ will be a port of the overlay graph of $\beta$ (with respect to $T$). Each such a port will either be non-redundant—in which case it will still be an edge of $S(T,\beta)$)—or it will be redundant—in which case it will be in the crossing sequence of some non-degenerate street and will be deleted to form $S(T,\beta)$ (left side of Figure 8). Therefore, for each street in $E_{S(T,\beta)}(\beta)$ (degenerate or not), we can use Part (c) of Lemma 2.3 and Part (b) of Theorem 3.2 to determine, in time $O(\|E_{T}(\alpha)\|\times m|T|)$, if there is an occurrence of an edge in $E_{T}(\alpha)$ in a street’s crossing sequence and delete from the complex $S(T,\beta)$ (the interior of) any such a street. Because the streets are disks in $\Sigma^{\prime}$, bounded by one or two segments of $\beta$, we note that, topologically, deleting the full interior of the streets that are transverse to some edge in $E_{T}(\alpha)$ is equivalent to puncturing a hole corresponding to a redundant port of the overlay graph by simply “enlarging” the hole (right side of Figure 8). This means that we have a cellular embedding of $\Sigma^{\prime\prime}$ which can be queried, in polynomial-time, for its topological type. Assuming that $\Sigma^{\prime\prime}$ has a component that is not a disk: we can use the techniques of [erickson2002optimally] (refer to Section 2.3 of [chambers2006splitting] for details) to find, in time $O(|T^{\prime}|)$ a non-contractable curve $s$ in $\Sigma^{\prime\prime}$ expressed as an SLP of complexity $O(m|T|+|T|)=O(m|T|)$. We finally compute $S(T,s)$, in time $O(m|T|^{2})$, to check whether a diagram is defined on each component of $\s\Sigma_{g}\backslash$ s. For such, we need only to find 1. the number of $\alpha$ and $\beta$ components on each side of $s$; and, 2. whether $s$ is not isotopic to an $\alpha$ or $\beta$ curve by looking for bounded cylinders, as described in the first case. ∎
@@
Given a normal curve $s$ in $T$ of complexity $\|N_{T}(s)\|=m^{\prime}$, one can find, in time $O(((m+m^{\prime})|T|)^{10})$, the diagram $(T,E_{T}(\alpha),N_{T}(\tau_{s}(\beta))$.
@@
This algorithm is described by [schaefer2008computing, Stefankovic]. If one requires only the intersection sequences $I_{T}(\tau_{s}(\beta))$, the algorithm can be executed in linear time on the inputs, but the output might just be standard with respect to the triangulation. Therefore, as in the proof of Theorem 3.5, we are forced to use the subroutines described by the same authors to cyclic reduce—either in time $O(((m+m^{\prime})|T|)^{10})$ with a deterministic algorithm or in time $O(((m+m^{\prime})|T|)^{3})$ with a randomized one—the multicurve given by $I_{T}(\tau_{s}(\beta))$. ∎
@@
A presentation of $\pi_{1}(M)$ can be computed in time $O(m|T|)$.